If a springis cut into two equal halfs then force constant of each half becomes 2k? How ? proof/explanation?

Grade:11

1 Answers

anil kumar
11 years ago

Force constant of a spring is inversely proportional to its length k¤1/l so k1*l1 = k2*l2 = kl(original spring) hence k1*l/2=k*l since length is halved hence k1=2k